OpenStudy (anonymous):

A new cruise ship line has just launched 3 new ships: the Pacific Paradise, the Caribbean Paradise, and the Mediterranean Paradise. The Caribean Paradise has 24 more deluxe staterooms than the Pacific Paradise. The Mediterranean Paradise has 29 fewer deluxe staterooms than twice the number of deluxe staterooms on the Pacific Paradise. Find the number of deluxe staterooms for each of the ships if the total number of deluxe staterooms for the three ships is 571.

hero (hero):

P + C + M = 571 C = P + 24 M = 2P - 29 P + (P + 24) + (2P - 29) = 571 4P - 5 = 571 4P = 576 P = 576/4 P = 144 C = 168 M = 259

OpenStudy (anonymous):

Lol, thanks Hero. I feel dumb forgetting that. Oh, nice little avatar.

hero (hero):

Yeah, it's just an avatar.

OpenStudy (anonymous):

Yeah, well thanks again. At least I haven't forgotten the y=mx+b formula you showed me a month ago.

hero (hero):

Oh goodie :D

OpenStudy (anonymous):

Oh just wait, I am getting ready for something harder.

hero (hero):

Of course you are. The next topic is always more advanced than the previous one.

OpenStudy (anonymous):

Especially when you never done it before.

hero (hero):

Yeah, you should cheat and learn the topics beforehand. That way, when the teacher introduces it to you, you'll already know it and she won't be able to surprise or embarrass you anymore.

OpenStudy (anonymous):

Polynomials is the next topic.

hero (hero):

How fun :D

OpenStudy (anonymous):

I will be hollering for you when I do it lol.

hero (hero):

But..but... you won't need me, cause you're going to learn it ahead of time :P

OpenStudy (anonymous):

Uh huh, if I can get past this part.

hero (hero):

I admit, it is kinda tricky in a way. Some people figure out how to do this without the algebra. Some others find that using the algebra is difficult. The rest just don't understand it at all.

hero (hero):

More intuitively, before learning algebra, I was able to figure out the answers using reasoning and trial and error. Now, I just use algebra. I'm pretty sure that after I learned algebra, my reasoning abilty waned greatly. Now I depend on the algebra a lot to answer such questions.

OpenStudy (anonymous):

Oh thanks.